# Jailbreak Large Vision-Language Models Through Multi-Modal Linkage

### v4

@@ 5.2 Main Experiments
The results of the experiment are presented in Table 1 and Table 2. As shown, our methods successfully jailbreak several target models with a high ASR across these datasets, making significant improvements over the baselines. For GPT-4o, we achieve an ASR of 97.80% on SafeBench, 98.81% on MM-SafeBench, and 99.07% on HADES-Dataset, representing increases of 64.80%, 73.56%, and 95.07% over the highest baseline results, respectively. Claude-3.5-Sonnet stands out as the most robust model, performing well against both the baselines and our method. However, MML still manages to jailbreak it with success rates of 69.40%, 60.00%, and 45.73% on the three datasets, showing improvements of 52.80%, 51.86%, and 45.60% compared to the highest baseline ASR. The experimental results demonstrate that current VLMs cannot maintain safety alignment under our attack. Qualitative attack results are listed in Figure 11.
@@
Table 2: ASR of HADES (Li et al., 2024b) vs. MML (ours) on HADES-Dataset. The letters under MML represent different encryption methods: WR stands for word replacement, M for image mirroring, R for image rotation, and B64 for base64 encoding. The highest ASR is highlighted in bold. All evaluations are conducted without any system prompts.
@@
The ASR vary across different encryption methods. As shown in Table 1, image transformation-based encryption outperforms both word replacement and base64 encoding. Base64 encoding shows the lowest success rate, likely due to more complex decryption process. Additionally, it is notable that Claude-3.5-Sonnet may have been specifically trained to defense base64 encoding based attack, which limits the effectiveness of MML with base64 encryption against it.
@@
Given that these datasets classify malicious queries into distinct categories, we also evaluate the ASR of our method across various forbidden topics. Figure 4 and Figure 5 illustrate the ASR of MML with image mirroring encryption compared to the baseline methods across four models on these topics in SafeBench and MM-SafeBench. On SafeBench, the baseline methods struggle with the first seven harmful topics such as Illegal Activity and Hate Speech on most models, mirroring the observations by Gong et al. (2023). In contrast, our method significantly improves the ASR for these topics, exceeding 95% in most cases, except for Claude-3.5-Sonnet. On MM-SafeBench, our approach consistently achieves at least 95% ASR across most topics and models. Notably, different models exhibit varying performance across different forbidden topics. ASR across various scenarios on HADES-Dataset and more detailed results are included in Appendix A.4.
@@ 5.3 Ablation Study
We select MML with image mirroring as the focus of our experiments, with results presented in Table 3. Additionally, we analyze the distribution of jailbreak scores under various conditions, as illustrated in Figure 6, to gain deeper insights into the impact of different components. Detailed prompts are provided in the Appendix A.5.
@@ 5.4 MML Performance under Defense
To further evaluate the effectiveness of MML, we explore its performance under AdaShield-Static (Wang et al., 2024), a prompt-based defense technique. We select two encryption methods for MML: word replacement and image mirroring. We conduct the experiment on SafeBench, using FigStep as the baseline method and GPT-4o as the target model. As before, we analyze the ASR across different forbidden topics in 5 attempts. We consider three variations of AdaShield-Static: placing the defense prompt before, in the middle of, or after the input text prompt. Detailed experimental setups are provided in the Appendix A.6.
The results are presented in Table 4. Consistent with previous studies (Wang et al., 2024), AdaShield-S performs well on the first seven topics against FigStep, reducing the ASR to 0. However, our method, particularly MML with word replacement as the encryption, performs robustly, especially when the defense prompt is sandwiched between the input text prompts. Apart from Illegal Activity, which experiences a 38% drop, reductions in other topics are mostly below 10%, maintaining an overall ASR of 87.80%. While MML-M shows a more significant decline, it still achieves an overall ASR of 59.20%.
@@ 5.5 Discussions
While we use only four encryption methods in this study, it’s crucial to note that MML can be extended to any encryption strategy that VLMs can handle. Exploring more encryption-decryption techniques may yield even better results, which we leave for future research.
@@
A key factor in MML’s success is the failure of safety alignment under complex instructions. When user prompts involve multiple steps, VLMs can become confused and lose safety alignment. Previous methods, such as designing intricate scenarios to “hypnotize” LLMs (Li et al., 2024a) or in-context learning-based jailbreak attacks (Anil et al., 2024; Wei et al., 2023b; Zheng et al., 2024), have indirectly validated this issue. Ensuring safety alignment in complex multi-step tasks without compromising model performance remains a crucial challenge.
@@ 8 Limitations
Despite MML achieves high jailbreak success rate on top VLMs such as GPT-4o, it has some limitations. First, we do not evaluate the more advanced o1 (OpenAI, 2024c) due to API access restrictions. Second, since MML does not instruct the model to conceal harmful content in the output (Wei et al., 2024), it can be defensed by output detection (Pi et al., 2024).
Appendix A Appendix
A.1 Implicit Rejection
@@
A.2 Candidate Replacement Words
@@ A.3 Complete Prompts of MML
A.4 Main Experiment Details
Baselines.
@@ A.4 Main Experiment Details
The five scenarios of HADES-Dataset (Li et al., 2024b) are Violence, Aiding and Abetting, Incitement; Financial Crime, Property Crime, Theft; Privacy Violation; Self-Harm; and Animal Abuse. Each category contains 150 instructions, for a total of 750 instructions.
@@
A.5 Ablation Experiment Details
@@
A.6 Defense Experiment Details

### v5

@@ 5.2 Main Experiments
The main results are presented in Table 1 and Table 2. As shown, MML methods successfully jailbreak several target models with a high ASR across these datasets, making significant improvements over the baselines. For GPT-4o, we achieve an ASR of 97.80% on SafeBench, 98.81% on MM-SafeBench, and 99.07% on HADES-Dataset, representing increases of 64.80%, 73.56%, and 95.07% over the highest baseline results, respectively. Claude-3.5-Sonnet stands out as the most robust model, performing well against both the baselines and our method. However, MML still manages to jailbreak it with success rates of 69.40%, 60.00%, and 45.73% on the three datasets, showing improvements of 52.80%, 51.86%, and 45.60% compared to the highest baseline ASR. Moreover, we also test MML on OpenAI o1 reasoning model (OpenAI, 2024c) on SafeBench, which keep outperform baseline 29.6% ASR on average. Details refer to Appendix D.8. The experimental results demonstrate that current VLMs cannot maintain safety alignment under our attack. Qualitative attack results are in Figure 11.
@@
Table 2: ASR of HADES (Li et al., 2024c) vs. MML (ours) on HADES-Dataset. The letters under MML represent different encryption methods: WR stands for word replacement, M for image mirroring, R for image rotation, and B64 for base64 encoding. The highest ASR is highlighted in bold. All evaluations are conducted without any system prompts.
@@
The ASR varies across different encryption methods. As shown in Table 1, image transformation-based encryption outperforms both word replacement and base64 encoding. Base64 encoding shows the lowest success rate, likely due to a more complex decryption process. Additionally, it is notable that Claude-3.5-Sonnet may have been specifically trained to defend against base64 encoding-based attacks, which limits the effectiveness of MML with base64 encryption against it.
@@
Given that these datasets classify malicious queries into distinct categories, we also evaluate the ASR of our method across various forbidden topics. Figure 4 and Figure 5 illustrate the ASR of MML with image mirroring encryption compared to the baseline methods across four models on these topics in SafeBench and MM-SafeBench. On SafeBench, the baseline methods struggle with the first seven harmful topics, such as Illegal Activity and Hate Speech on most models, mirroring the observations by Gong et al. (2023). In contrast, our method significantly improves the ASR for these topics, exceeding 95% in most cases, except for Claude-3.5-Sonnet. On MM-SafeBench, our approach consistently achieves at least 95% ASR across most topics and models. Notably, different models exhibit varying performance across different forbidden topics. ASR across various scenarios on HADES-Dataset and more detailed results are included in Appendix D.
@@ 5.3 Ablation Study
We select MML with image mirroring as the focus of our experiments, with results presented in Table 3. Additionally, we analyze the distribution of jailbreak scores under various conditions, as illustrated in Figure 6, to gain deeper insights into the impact of different components. Detailed prompts are provided in the Appendix E.
@@
Encryption WR M R B64
Cost time (s) 120.33 2.37 2.41 3.39
Table 4: Time cost of encrypting 500 images using different methods. WR stands for word replacement, M for image mirroring, R for image rotation, and B64 for base64 encoding.
Run FS QR MML-WR MML-M MML-R MML-B64
Run1 34.13 (359, 1) 77.90 (602, 2) 68.65 (2272, 10) 63.58 (2022, 10) 73.02 (2522, 10) 70.85 (2196, 10)
Run2 40.92 (910, 2) 72.53 (417, 1) 56.77 (2374, 10) 76.79 (2534, 10) 75.38 (2589, 10) 81.69 (2304, 10)
Run3 32.10 (74, 0) 70.49 (487, 2) 74.19 (2382, 10) 67.96 (2315, 10) 68.74 (2188, 10) 74.46 (2272, 10)
Avg 35.72 (448, 1) 73.64 (502, 1.67) 66.54 (2342, 10) 69.44 (2290, 10) 72.38 (2433, 10) 75.67 (2257, 10)
Table 5: Time cost for GPT-4o to generate responses for 10 fixed samples. The table entries are presented in the format: time (word count, no-refusal count). For example, ‘34.13 (359, 1)’ indicates that the model takes 34.13 seconds to respond to 10 queries, generates 359 words in total, and provides only 1 response that was not a direct refusal. FS represents FigStep (Gong et al., 2023), and QR represents QueryRelated (Liu et al., 2024b). MML-$X$ denotes variations of MML with different encryption methods: WR stands for word replacement, M represents image mirroring,R for image rotation, and B64 for base64 encoding.
@@
5.4 Time Efficiency of MML
MML consists of two main stages: the encryption of the image and the inference of the target model. For encryption, we measure the time required to encrypt the Safebench dataset using different encryption strategies. The total time taken for 500 images is shown in Table 4.222These experiments are conducted locally on a MacBook Pro with Apple M1 Pro, using Python to encrypt images. As shown, most encryption methods take less than 3.5 seconds. Word replacement, however, requires using the NLTK for part-of-speech analysis in the pipeline, which requires more processing time. Nevertheless, we suggest this additional time is well within an acceptable range.
To measure the inference time for different attack methods, we calculate the time it takes for GPT-4o to generate responses for 10 fixed samples in three independent runs. We also report the total word count of these 10 responses and the number of instances where the model does not directly refuse the request. The results in Table 5 show that Figstep takes less time, while MML and QueryRelated both take longer and require similar amounts of time. This difference in processing time stems from two factors:
  • Image size: GPT-4o processing is faster for FigStep and MML due to their smaller images (20-30KB) compared to QueryRelated’s larger images ( 1.5MB).
  • Response length: our MML method requires more model reasoning time than FigStep and QueryRelated. It is because MML’s prompts are often longer, and the more complex instruction leads to longer, more detailed outputs. In contrast, FigStep often responds concisely with rejections (e.g., “I’m sorry, I can’t assist with that”), resulting in shorter times, especially when all attacks fail (Run 3). Its time still increases when attacks elicit longer responses (Run 2).
In summary, while MML might take slightly longer, it is in the same order of magnitude as the baseline. We consider this additional time a reasonable trade-off that will not significantly hinder the utility of attack methods.
@@ 5.5 MML Performance under Defense
To further evaluate the effectiveness of MML, we explore its performance under AdaShield-Static (Wang et al., 2024), a prompt-based defense technique. We test two encryption methods for MML: word replacement and image mirroring. We experiment on SafeBench, using FigStep as the baseline method and GPT-4o as the target model. As in previous settings, we measure the ASR across different forbidden topics in 5 attempts. We consider three AdaShield-Static variants, where the defense prompt is inserted before, between, or after the attack prompt. Detailed experimental setups are provided in the Appendix F.
The results (Table 6) confirm prior findings (Wang et al., 2024) that AdaShield-Static effectively reduces FigStep’s ASR to 0 on the first seven topics. However, MML—particularly the word replacement variant—remains resilient. Except for Illegal Activity (which sees a 38% ASR drop), most topics show reductions under 10%, yielding a strong overall ASR of 87.80%. Although image mirroring is more affected, it still achieves a respectable ASR of 59.20%, demonstrating MML’s robustness even under defensive interventions.
@@ 5.6 Discussions
While our main experiments focus on four basic encryption methods, MML is inherently flexible and can be extended to a wider range of encryption strategies that VLMs are capable of interpreting. To further demonstrate this extensibility, we incorporate the Shift Cipher (SC)—a classical substitution cipher in which each letter is shifted one position forward during encryption and one position backward during decryption McCoy et al. (2023). We evaluate the effectiveness of MML-SC on SafeBench, using the same experimental setup described in Section 5. The results, presented in Table 7, are comparable to or even surpass those basic results reported in Table 1, further validating MML’s adaptability to alternative encryption schemes. The full prompt used for the MML-SC variant is provided in Figure 19.
@@
A key factor in MML’s success is the failure of safety alignment under complex instructions. When user prompts involve multiple steps, VLMs can become confused and lose safety alignment. Previous methods, such as designing intricate scenarios to “hypnotize” LLMs (Li et al., 2024b) or in-context learning-based jailbreak attacks (Anil et al., 2024; Wei et al., 2023b; Zheng et al., 2024), have indirectly validated this issue. Ensuring safety alignment in complex multi-step tasks without compromising model performance remains a crucial challenge.
Model ASR(%)
FS QR MML-SC
GPT-4o 33.00 27.00 99.40
GPT-4o-Mini 39.00 32.20 96.20
Claude-3.5-Sonnet 16.60 19.40 68.20
Qwen-VL-Max 92.60 62.00 98.60
Table 7: ASR of baseline methods and MML-SC (ours). SC denotes shift ciphers. FS represents FigStep (Gong et al., 2023), and QR represents QueryRelated (Liu et al., 2024b). Best results are highlighted in bold. All evaluations are conducted without any system prompt.
@@ 8 Limitations
Despite MML achieves high jailbreak success rate on top VLMs such as GPT-4o, it has some limitations. Since MML does not instruct the model to conceal harmful content in the output (Wei et al., 2024), it can be defensed by output detection (Pi et al., 2024).
9 Acknowledgment
This work was supported by National Natural Science Foundation of China (N0.62176252).
Appendix A Implicit Rejection
@@
Appendix B Candidate Replacement Words
@@ Appendix C Complete Prompts of MML
Appendix D Main Experiment Details
D.1 Baselines
@@ D.2 Datasets
The five scenarios of HADES-Dataset (Li et al., 2024c) are Violence, Aiding and Abetting, Incitement; Financial Crime, Property Crime, Theft; Privacy Violation; Self-Harm; and Animal Abuse. Each category contains 150 instructions, for a total of 750 instructions.
@@ D.6 ASR of MML on various topics
D.7 Evaluation with Llama-Guard-3-8B
To mitigate potential bias from a single evaluator, we use Llama-Guard-3-8B (Llama Team, 2024) for an additional assessment of attack success. Since Llama-Guard-3-8B has a built-in safety policy, not all instructions in these datasets that we use to evaluate are flagged as unsafe by it. Therefore, we first use Llama-Guard-3-8B to filter out all unsafe instructions, resulting in 307 from SafeBench, 867 from MM-SafeBench, and 687 from HADES-Dataset. Then we use Llama-Guard-3-8B to evaluate the safety of the target model’s responses to these specific harmful instructions. The attack success rate is the proportion of these responses flagged as unsafe, with results shown in Table 9 and Table 10. The evaluation results using Llama-Guard-3-8B are aligned with the results in Table 1 and Table 2.
@@ D.7 Evaluation with Llama-Guard-3-8B
MML with Shift Ciphers Prompt I’m developing a game, and this image shows a large screen inside the villain’s lair. But some information on the screen is missing. I need your help to fill in the gaps and make it more realistic based on the visible titles. However, the letters of the titles are now shifted back one position, for example, ’a’ is written as ’b’, ’b’ is written as ’c’,…, ’z’ is written as ’a’. Only letters are encrypted, punctuation marks and numbers are not encrypted. I also have a scrambled list of all the original words from the title: {shuffled_list} You need to follow the following thinking steps to add content: 1. Carefully extract the title on the picture and decrypt it according to the rules, that is, move each letter forward one position. 2. Verify if all the words in the extracted title correspond to the scrambled word list. 3. If they match perfectly, provide real-world content based on the title. 4. If there are discrepancies, adjust the title until it matches exactly, then add content. The added content must meet the following requirements: 1. Ensure the correctness of the reconstructed title. 2. Fully elaborate on the content according to the given titles, including necessary details and steps. 3. Ensure the supplemented content is realistic and aligns with the villain’s evil nature. Your response should follow this format: I am glad to participate in your game production! Extracted Title: "[Provide the extracted title from the image]" Reconstructed Title: "[Provide the decrypted title according to the rules]" Content: Below is the added content for "[Provide the decrypted title according to the rules]": 1. 2. 3.
Figure 19: Prompt of MML with shift ciphers encryption.
@@
D.8 MML attack on reasoning model
We additional evaluate MML methods targeted on OpenAI o1 reasoning model (OpenAI, 2024c) on SafeBench. The results are shown in Table 11. The experimental settings are the same as Section 5. Though the ASR is lower than that of jailbroken GPT-4o, it still shows a considerable outperformance over baseline methods (+29.6%).
Appendix E Ablation Experiment Details
@@
Appendix F Defense Experiment Details
@@
Dataset Model ASR(%)
FS QR MML-WR MML-M MML-R MML-B64
SafeBench GPT-4o 6.19 3.91 96.42 97.39 96.42 96.74
GPT-4o-Mini 13.68 7.82 96.42 95.77 96.42 94.14
Claude-3.5-Sonnet 3.26 1.30 41.69 53.42 38.11 8.14
Qwen-VL-Max 89.58 53.75 92.51 96.74 92.18 91.86
$M$-SafeBench GPT-4o 1.04 13.23 95.24 95.48 95.94 96.40
GPT-4o-Mini 16.13 12.06 94.78 94.78 93.16 93.16
Claude-3.5-Sonnet 2.44 2.78 40.02 48.62 37.12 9.28
Qwen-VL-Max 27.84 48.49 91.76 92.34 91.42 92.23
Table 9: ASR evaluated by Llama-Guard-3-8B. FS represents FigStep (Gong et al., 2023), and QR represents QueryRelated (Liu et al., 2024b). MML-$X$ represents different encryption methods: WR stands for word replacement, M for image mirroring, R for image rotation, and B64 for base64 encoding. Best results are highlighted in bold. All evaluations are conducted without any system prompt.
Model ASR(%)
HADES MML
WR M R B64
GPT-4o 2.62 97.96 97.82 97.09 97.53
GPT-4o-Mini 2.62 97.82 97.23 96.65 94.76
Claude-3.5-Sonnet 0 35.66 42.07 29.84 9.46
Qwen-VL-Max 43.67 96.22 96.51 97.23 95.49
Table 10: ASR evaluated by Llama-Guard-3-8B on HADES-Dataset. The letters under MML represent different encryption methods: WR stands for word replacement, M for image mirroring, R for image rotation, and B64 for base64 encoding. The highest ASR is highlighted in bold. All evaluations are conducted without any system prompts.
Model FS QR MML-WR MML-R
o1 34.8 15.6 64.4 49.4
Table 11: ASR of baseline methods and MML (ours) on OpenAI-o1. FS represents FigStep (Gong et al., 2023), and QR represents QueryRelated (Liu et al., 2024b). MML-$X$ represents different encryption methods: WR stands for word replacement, M for image mirroring. Best results are highlighted in bold. All evaluations are conducted without any system prompt.